OpenStudy (anonymous):

if tan(x)=-5/12 then sin(x)=?,cos(x)=?

OpenStudy (johnweldon1993):



OpenStudy (johnweldon1993):

We know tangent = opposite/adjacent.. So you make a triangle....pick an angle...and use tanx = -5/12 to place the numbers as I did

OpenStudy (johnweldon1993):

And since you have 2 legs of the triangle and you want the hypotenuse you can use the Pythagorean Theorem to solve for that hypotenuse...we found to be 13

OpenStudy (johnweldon1993):

So now... sinx = opposite/hypotenuse... From that angle 'x' we look at the opposite side to it (-5) and the hypotenuse (13) so sinx = -5/13

OpenStudy (johnweldon1993):

And cosx same thing... cos = adjacent/hypotenuse From that angle we want the adjacent side (12) and the hypotenuse (13) so cosx = 12/13

OpenStudy (anonymous):

i understand the process.thanks